OpenStudy (zzr0ck3r):

G is a group for a in G define C_a = { x in G s.t. xa = ax} prove that there is a one-to-one correspondence between the set of all the conjugates of a and the set of all the cosets of C_a

OpenStudy (zzr0ck3r):

hint: use the fact that x^-1*a*x = y^-1*a*y iff C_a*x = C_b*y

terenzreignz (terenzreignz):

What's a conjugate?

OpenStudy (zzr0ck3r):

for a in G its the for x^-1*a*x for x in G

OpenStudy (zzr0ck3r):

form*

OpenStudy (kinggeorge):

This makes intuitive sense, but I'm not immediately sure of the best way to approach it.

OpenStudy (zzr0ck3r):

do I need to show they are the say size?

OpenStudy (zzr0ck3r):

same size*

OpenStudy (kinggeorge):

That would be one way. To show that there are the same number of conjugates as there are cosets.

terenzreignz (terenzreignz):

Wait... conjugates is \[\Large \left\{a\in G \quad \left|\quad xax^{-1} = a \quad \forall x \in G\right.\right\}\]

terenzreignz (terenzreignz):

^was a question

OpenStudy (zzr0ck3r):

if a is in G, a conjugate of a is any element of the form x*a*x^-1 where x is in G.

OpenStudy (zzr0ck3r):

so not all x

terenzreignz (terenzreignz):

ahh, okay understood, sorry :D

OpenStudy (kinggeorge):

Conjugates of \(g\in G\) are \[\{a\in G|hgh^{-1}=a \text{ for some }h\in G\}\]

OpenStudy (zzr0ck3r):

much nicer:)

OpenStudy (zzr0ck3r):

number of cosets will be ord(g)/ord(C_a) I think?

OpenStudy (kinggeorge):

I think so.

terenzreignz (terenzreignz):

So we just have to find an injective map?

OpenStudy (kinggeorge):

Bijective map.

terenzreignz (terenzreignz):

It did just say one-to-one...

OpenStudy (zzr0ck3r):

1-1 correspondence means both

terenzreignz (terenzreignz):

ohh. Okay.

OpenStudy (zzr0ck3r):

its silly wording

OpenStudy (zzr0ck3r):

well not silly but ...

OpenStudy (anonymous):

That hint really does the whole problem for you.

OpenStudy (kinggeorge):

agreed^

OpenStudy (zzr0ck3r):

I guess I just don't know what im trying to show. would it be the size?

terenzreignz (terenzreignz):

Maybe start with defining a set \[\Large G_a = \left\{xax^{-1} \quad \left|\quad x\in G\right.\right\}\]

OpenStudy (anonymous):

If A is the set of all conjugates of a, then create the map:\[f: A\rightarrow G\big/C_a\]\[f(x^{-1}ax)=C_ax\]

OpenStudy (anonymous):

Or use teren's notation, thats much better. (instead of my A).

OpenStudy (zzr0ck3r):

ok so this is the map and I show that it is 1-1 and onto?

terenzreignz (terenzreignz):

And now, all that's left is to show that said mapping is both injective and surjective :D

terenzreignz (terenzreignz):

(I still don't know how to make that happen, but it's there)

OpenStudy (anonymous):

right. and that hint does all of it except for the onto, which is easy enough.

OpenStudy (anonymous):

You also need to show that the map is well defined. (which again the hint shows).

terenzreignz (terenzreignz):

Showing it's onto may have been simpler than I thought... Define an arbitrary coset of \(\large C_a\) and show that there is a conjugate of a that will be mapped to said coset.

OpenStudy (zzr0ck3r):

so if I take and conjugate xax^-1 and another yay^-1 and assume f(one) = f(the other) C_a*x = C_a*y iff our two elements of equal. thus its 1-1 to show onto take y in G/C_a and this is where im a little lost

OpenStudy (zzr0ck3r):

I need to show that I can obtain y from some f(b) where b is a conjugate of a

OpenStudy (anonymous):

So given \[C_ay\]with \[y\in G\]can you think of a conjugate of a that maps to it?

OpenStudy (anonymous):

\[f(\_a\_)=C_ay\]what should go in the blank?

OpenStudy (zzr0ck3r):

sorry my wife walked in the door

OpenStudy (zzr0ck3r):

yay^-1

terenzreignz (terenzreignz):

yay :)

OpenStudy (anonymous):

right.

OpenStudy (zzr0ck3r):

ty sorry about that

OpenStudy (anonymous):

lol, no worries.

OpenStudy (zzr0ck3r):

and then as far as it being well defined I just need to show that ever element has an image I can do that:)

terenzreignz (terenzreignz):

More like you have to show that each element (in the domain) will have only one image.

OpenStudy (anonymous):

Good :) its important to show the function is well defined, you might want to put that before you show the 1-1 and onto.

OpenStudy (zzr0ck3r):

right

OpenStudy (zzr0ck3r):

@joemath314159 this might be silly but.... the way you have the function f(x^-1*a*x) but the book defines conjugates as x*a*x^-1 does this matter? also I am a little confused on how to show that it is well defined. what exactly do I need to show?

OpenStudy (anonymous):

in the big picture, there is no difference in\[x^{-1}ax\]versus \[xax^{-1}\]The only thing that will change are the details, but as long as you stick with only one, the main results will look exactly the same. I would choose whichever way your book/class/professor defines conjugation.

OpenStudy (zzr0ck3r):

well the book says, when it defines conjugates x*a*x^-1 but when we use that hint the book says x^-1*a*x = y^-1*a*y iff C_ax = C_ay

OpenStudy (zzr0ck3r):

but we use the fact that our element is from the set of conjugates so I don't know how to show 1-1

OpenStudy (anonymous):

As far as a function being well defined, in general what you are trying to show is\[x=y\Longrightarrow f(x)=f(y)\]That statement looks trivial (and is in most cases), but in our setting, where we defined:\[f(x^{-1}ax)=C_ax\]we defined the output of the function based on on particular representation of the conjugate. If x is not equal to y and:\[x^{-1}ax=y^{-1}ay\]Then its clear that:\[f(x^{-1}ax)=f(y^{-1}ay)\]but:\[f(x^{-1}ax)=C_ax,f(y^{-1}ay)=C_ay\]Without the hint, we really dont know for sure that:\[C_ax=C_ay\]. If they didnt, our function would not be well defined.

OpenStudy (anonymous):

and to further highlight that the way you define a conjugate really doesn't change much, note that if we take \[x^{-1}ax\]to be a conjugate, let:\[g=x^{-1}\]Then you get:\[gag^{-1}\]The only difference is who you view as the inverse, the one on the left, or the one on the right.

OpenStudy (zzr0ck3r):

ok that's what I was thinking, but didn't know how to say it:)